# R2C2-Coder: Enhancing and Benchmarking Real-world Repository-level Code Completion Abilities of Code Large Language Models

### v1

@@ B.1 Details on Abstract Context Extraction
Figure 8: A case found in CrossCodeEval [16]. As Java supports the package statement, developers could call cross-file APIs in the same package. Such kinds of implicit dependencies are ignored by Ding et al. [16].
@@ B.2 Quality Control Procedure
We first define the ground-truth completion result as the “reference”, and design a series of rule-based and model-based post-processing filters to ensure the quality of the dataset as follows. First, we filter examples if (1) too short ($<10$ tokens) reference. (2). too long ($5$ lines) reference. Second, we keep that the percentage of the multiple lines for the reference is larger than 30%. Moreover, to ensure that the reference isn’t predictably inferred solely from the current file (possibly owing to strong clues in function names and comments), we feed the examples (input prompts) to DeepSeekCoder-1.3B model [21] to complete the statement and remove the exact matches. This step removes 57.1%, 39.8%, 48.0%, and 47.2% of the samples for Java, Python, TypeScript, and C#, respectively. Finally, we perform human annotations on a subsample of R2C2-Bench and found that the dataset has a satisfactory quality to serve the goal of cross-file code completion.
@@ C.1 Details of the baseline models
StarCoder [32] is a generative decoder-only LM series whose size varies from 1B to 15.5B. Trained on the Stack [30] dataset, StarCoder supports a maximum context length of 8K.
DeepSeekCoder [21] is a suite of code models with sizes ranging from 1.3B to 33B trained on a 2-trillion-token manual collected corpus. Benefit from FIM strategy [8] and Rotary Position Embedding (RoPE) [52], these models are capable of efficient code generation and infilling under a 16K context window.
Code Llama [49] is a family of large language models based on Llama 2 [55] with 7B, 13B, 34B, and 70B parameters. These models are trained on sequences of 16K tokens and capable of inference on 100K. We conduct experiments with the foundational Code Llama models, not the Python-specialized versions or instruction-tuned ones.
@@ C.2 Details of CrossCodeEval and CrossCodeEval+
CrossCodeEval [16] is an arising benchmark for cross-file code completion that is built on a diverse set of real-world, open-sourced, permissively-licensed Github repositories in four popular programming languages: Python, Java, TypeScript, and C#. To create accurate and high-quality test samples, the authors adopt static analysis and filter out those samples solved by starcoderbase-1B within in-file only context.
CrossCodeEval+. We collect the original code bases of CrossCodeEval [16] then parse and extract the cross-file dependencies with our R2C2-Enhance. During retrieval, the abstracts and the snippets are placed together in the candidate pool (Sec. 3.1.1). Note that we keep the cursor positions and the expected outputs the same as those in CrossCodeEval. Here, we name the reconstructed CrossCodeEval as CrossCodeEval+.
C.3 Execution Time Benchmarking

### v3

@@ B.1 Details on Abstract Context Extraction
Figure 8: A case found in CrossCodeEval (Ding et al., 2023). As Java supports the package statement, developers could call cross-file APIs in the same package. Such kinds of implicit dependencies are ignored by Ding et al. (2023).
@@ B.2 Quality Control Procedure
We begin by establishing the ground-truth completion result as the “reference” and create a set of post-processing filters (i.e., rule-based and model-based rules) to enhance dataset quality. Initially, we eliminate examples based on two criteria: (1) references that are too short (fewer than 10 tokens) and (2) those that are overly long (exceeding 5 lines). Next, we ensure that more than 30% of our references are in multiple lines. To prevent predictable inference from the current file—potentially due to prominent indicators in function names and comments—we input the examples (input prompts) into the DeepSeekCoder-1.3B model (Guo et al., 2024a) for completion and discard any exact matches. This step results in the removal of 57.1%, 39.8%, 48.0%, and 47.2% of samples for Java, Python, TypeScript, and C#, respectively. Lastly, we conducted human annotations on a subset of R2C2-Bench and observed that the dataset possesses sufficient quality for cross-file code completion purposes.
@@ C.1 Details of the baseline models
StarCoder (Li et al., 2023) is a generative decoder-only LM series whose size varies from 1B to 15.5B. Trained on the Stack (Kocetkov et al., 2022) dataset, StarCoder supports a maximum context length of 8K.
DeepSeekCoder (Guo et al., 2024a) is a suite of code models with sizes ranging from 1.3B to 33B trained on a 2-trillion-token manual collected corpus. Benefit from FIM strategy (Bavarian et al., 2022) and Rotary Position Embedding (RoPE) (Su et al., 2024), these models are capable of efficient code generation and infilling under a 16K context window.
Code Llama (Roziere et al., 2023) is a family of large language models based on Llama 2 (Touvron et al., 2023) with 7B, 13B, 34B, and 70B parameters. These models are trained on sequences of 16K tokens and capable of inference on 100K. We conduct experiments with the foundational Code Llama models, not the Python-specialized versions or instruction-tuned ones.
@@ C.2 Details of CrossCodeEval and CrossCodeEval+
CrossCodeEval (Ding et al., 2023) is an arising benchmark for cross-file code completion that is built on a diverse set of real-world, open-sourced, permissively-licensed Github repositories in four popular programming languages: Python, Java, TypeScript, and C#. To create accurate and high-quality test samples, the authors adopt static analysis and filter out those samples solved by starcoderbase-1B within in-file only context.
CrossCodeEval+. We collect the original code bases of CrossCodeEval (Ding et al., 2023) then parse and extract the cross-file dependencies with our R2C2-Enhance. During retrieval, the abstracts and the snippets are placed together in the candidate pool (Sec. 3.1.1). Note that we keep the cursor positions and the expected outputs the same as those in CrossCodeEval. Here, we name the reconstructed CrossCodeEval as CrossCodeEval+.
(a) StarCoder
(b) CodeLlama
(c) DeepSeekCoder
Figure 12: Performance on CrossCodeEval after R2C2-Enhanced tuning with various input lengths.
C.3 Scalability of Model Performance on Various Input Lengths
As depicted in Figure 12, we report results produced by code models on CrossCodeEval (Ding et al., 2023) when the input lengths of R2C2-Enhanced tuning range in {512, 1024, 2048, 4096} tokens. There exists a scaling law that the performance is improved when input length increases.
C.4 Execution Time Benchmarking